OpenStudy (trisha001):

The length of a rectangle is 6 meters more than its width. If the area of the rectangle is 135 square meters, find its dimensions.

OpenStudy (trisha001):

wait so is it x+6 or x2+6x

OpenStudy (trisha001):

I'm a little confused sorry

Mehek (mehek14):

hold on

Mehek (mehek14):

\(2L+2W=135\\L=6+W\\2(6+W)+2W=135\)

OpenStudy (anonymous):

w x w + 6 = 135 w2 + 6w - 135 =0 Ie (w - 9)(w + 15) = 0 width = 9m (length = 15m)

Mehek (mehek14):

now distribute \(2(6+W)\)

OpenStudy (trisha001):

Now I'm really confused. I don't understand this at all

Mehek (mehek14):

do you know distribution?

OpenStudy (trisha001):

Not really. Could you explain please?

Mehek (mehek14):

\(\Large{a(b+c)=a*b+a*c}\) Multiply what's on the outside with everything on the inside.

Mehek (mehek14):

@trisha001 can you do it now?

OpenStudy (trisha001):

so I solve a*b+a*c first? and multiply it with a(b+c) ?

Mehek (mehek14):

no you multiply a(b+c) which gives you a*b + a*c do that with 2(6+W)

OpenStudy (trisha001):

so now I multiply 2(6+w) a(b+c) or 2(6+w) and a*b+a*c

OpenStudy (trisha001):

I'm sorry for being so confused

Mehek (mehek14):

a(b+c) was just an example '-' 2(6+W) = 2*6 + 2*W

Mehek (mehek14):

@Serenity74

OpenStudy (trisha001):

alright sorry

OpenStudy (trisha001):

when I did the problem I got 0=0 is that correct?

Mehek (mehek14):

???

OpenStudy (trisha001):

2(6+W) = 2*6 + 2*W

Mehek (mehek14):

2*6 = ?

OpenStudy (trisha001):

I can't figure it out. So thank you for taking your time to help but I'm having a lot of trouble so it's okay.

Mehek (mehek14):

Oh oke. . . Well I have to go off. . . So I'll see if someone else can help :)

OpenStudy (trisha001):

I finally got it now so thank you anyways but I figured it out ((: I really do appreciate it though.

Mehek (mehek14):

yw :)